# Real-Time Position-Aware View Synthesis from Single-View Input

### v1

@@ 3.1. Positional Embedding
We define the target position by the coordinate vector , where , and represent translations associated with angular coordinates. Given limited baseline shifts of LF reconstruction, directly using normalized values of can suffice for training the network on discrete locations but querying for any continuous values of .
Coordinate Normalization: The target coordinates are normalized into to ensure consistency across varying ranges of , . The normalization process is given by
This normalization step ensures that the coordinates lie within a fixed range, which is crucial for subsequent encoding.
Input Channel: To incorporate the normalized coordinates into the network, the values of are expanded into a tensor , where the first channel holds the values, and the second channel holds the values
This tensor has the same spatial resolution as the input image, ensuring that positional information is aligned with the image features spatially.
@@ 3.2. Rendering Network
The role of the rendering network in PLFNet is to synthesize a novel target image from the input image and the positional embedding , obtained from the positional embedding stage. Our network architecture is inspired by LFSphereNet (Gond et al., 2023) and consists of two distinct encoder branches followed by a decoder network. These branches work in tandem to extract and merge spatial and positional features, ultimately generating high quality novel views.
@@ 3.2.2. Encoder II - Position-Aware Feature Extractor
The second encoder, , integrates both the input image and the positional embedding to adaptively capture position-specific features. Unlike the first encoder, which only processes the image, focuses on refining the feature maps based on the target position.
@@
where is the number of channels in . In our case, for LF Reconstruction.
This stacked tensor is passed through a series of convolutional layers that preserve the spatial dimensions, ensuring pixel-level positional information is retained throughout the encoding process, as demonstrated by LFSphereNet (Gond et al., 2023). The encoder produces a feature map , where is the number of feature channels, and skip connections are generated from intermediate layers for later use in the decoder.
@@ 3.3. Training Loss
To optimize our PLFNet, we employ a combination of the L1 loss and the Multi-Scale Structural Similarity (MS-SSIM) (Wang et al., 2003) loss. This combination leverages the complementary strengths of each loss function. The L1 loss minimizes pixel-wise differences between the predicted sub-aperture images and the ground truth images , which helps to maintain accurate reconstruction of intensity values. The MS-SSIM loss, on the other hand, is a perceptually motivated metric that focuses on preserving structural information and texture across multiple scales, which correlates better with human visual perception as shown in image restoration task (Zhao et al., 2016).
The total loss function is defined as a weighted sum of the L1 and MS-SSIM losses
where is a weighting factor that balances the contribution of the two losses. This combined loss function enables the network to generate high-quality light field reconstructions that are both pixel-accurate and perceptually consistent.
3.4. Implementation: PLFNet
The input of PLFNet is a single RGB image and the target position . Here and which are similar to the spatial resolution of input image . We set because the datasets (Srinivasan et al., 2017; Raj et al., 2016; Rerabek and Ebrahimi, 2016) are organized as a 2D grid, with only two angular coordinates considered. PLFNet was implemented in PyTorch and trained with the following hyperparameters: batch size of 24, learning rate of 0.003, adam optimizer, MS-SSIM and L1 loss. After each 30 epochs, the learning rate was decreased by a factor of 0.2. The network was trained for 150 epochs on 2 GPUs (Nvidia A40) within a computing cluster with Intel Xeon Gold 6338 CPUs. The training duration was 48 hrs. For the dataset split, 75% of the data was used for training, 12.5% for validation and 12.5% for testing.
4. Method: View Synthesis - PLFNet+
Our approach for view synthesis builds upon PLFNet by extending the positional embedding to handle wider baselines with translation along the -axis. Here, represents the synthesized image corresponding to the target viewpoint. Our approach consists of two key stages as shown in Figure 2: first, the target position is embedded into a high-dimensional space through positional embedding (discussed in Section 4.1), and second, novel views are rendered from these positions, which is detailed in Section 4.2.
4.1. Extended Positional Embedding
We define the target position by the coordinate vector , where , , and represent translations from the input image . As the baseline increases and the search space extends along the axis, it becomes necessary to project the target coordinates into a higher-dimensional space to accurately capture positional information. To address this, we employ a positional embedding strategy that combines positional encoding, inspired by NeRF (Mildenhall et al., 2021), along with a MLP for enhanced representation of the target coordinate .
Coordinate Normalization: Similar to the normalization strategy in Section 3.1, we convert the target coordinates into their normalized form .
where is the normalized 3D coordinate vector, represents the encoding depth, and . The parameter is a scaling constant based on the domain of . For our experiments, we use , therefore yields an output of size dimensions, which provides enhanced positional awareness (see Table 6 for ablation study results).
Learning Based Reprojection: Once the positional encoding has been computed, we further enhance the representation by passing it through an MLP. This MLP maps the low-dimensional encoded vector into a higher-dimensional space, and it has proven to work well with scene regression task as shown in NeRF (Mildenhall et al., 2021). This process allows for a more expressive and flexible feature space, and our approach is similar to the adaptive layer normalization techniques used in NViST (Jang and Agapito, 2024).
Let be the output of the positional encoding, where is the dimensionality of the encoding. The MLP projects this vector into a high-dimensional space , where is chosen based on the desired resolution. Mathematically, the projection is represented as
After this, the vector is reshaped into a feature map of size , where . This reshaping operation can be expressed using the inverse vectorization notation as
Finally, the feature map is upsampled to match the spatial resolution of the input image, denoted by . This upsampling is performed using a scaling factor , as follows
4.2. Rendering Network
The rendering network remains similar to Section 3.2. However, in the case of view synthesis, the positional embedding is denoted by a single channel, simplifying the input to the second encoder as
The output of the rendering network is the target image .
4.3. Training Loss
For the view synthesis task, we extend the loss function previously defined in Section 3.3 to better capture fine-grained details, especially with larger baselines. We maintain the same combination of L1 loss and MS-SSIM loss, which focuses on minimizing pixel-wise differences and preserving structural consistency, respectively. However, to further enhance the quality of the synthesized views, particularly in the frequency domain, we incorporate the Focal Frequency Loss (FFL) (Jiang et al., 2021) into the total loss function.
@@ 4.3. Training Loss
The FFL addresses the spectral bias in neural networks by adaptively focusing on the frequency components that are hardest to synthesize. This allows the model to improve the reconstruction of high-frequency details, which are critical for realistic novel view generation.
The FFL is defined as
@@
where and represent the real and predicted frequency components at position in the Fourier domain, and is a dynamic spectrum weight that emphasizes hard-to-synthesize frequencies.
Thus, the total loss function for PLFNet+ is updated to
where is the Focal Frequency Loss, and is a weight balancing its contribution relative to the L1 and MS-SSIM losses.
This augmented loss function allows PLFNet+ to generate sharper, more realistic novel views by not only optimizing for spatial accuracy but also addressing frequency domain discrepancies.
4.4. Implementation: PLFNet+
We keep the same implementation as in PLFNet (described in Section 3.4, but set , . We also set as the positional maps is single channel. Due to limitation of training data when using our synthetic dataset we employ multi stage training strategies to reach satisfactory results in training process.
@@ 4.4.1. Multi Stage Training
Stage 1 serves as a preconditioning step, training the network on a large-scale light field dataset (Srinivasan et al., 2017). While the baseline between the views in each light field is relatively narrow, the dataset’s large scale enables the network to learn both high- and low-level image features, implicitly understanding depth. In this stage, the network is conditioned to synthesize novel views within a constrained light field volume, similar to the reconstruction process described in PLFNet (Section 3). Once the network demonstrates proficiency in this task, we transition to the next stage.
@@
Stage 3 involves training the network on a synthetic dataset containing random camera viewpoints within an cubical volume. The increase in the spatial search space necessitates the use of the positional embedding module for accurate view synthesis. In this stage, the network is fine-tuned to handle more complex camera translations and larger scene variations, fully utilizing the positional embedding to enhance rendering precision.
@@ 5. Experimental Evaluation
In this section, we present a comprehensive evaluation of our proposed PLFNet and PLFNet+ architectures. We begin by assessing the performance of PLFNet+ for view synthesis on datasets with varying baselines and complexities. Following this, we demonstrate the capability of PLFNet in reconstructing full light fields from a single input image. The experiments for view synthesis are detailed in Section 5.1, while light field reconstruction results are discussed in Section 5.2.
5.1. View Synthesis
@@
| PLFNet+ (Ours) | 28.8023 | 0.9003 | 0.9655 | 0.7516 | 0.1586 | 0.1007 | |
@@
| PLFNet+ (Ours) | 26.3339 | 0.8334 | 0.9307 | 0.5633 | 0.2191 | 0.1947 | |
| SinMPI (Pu et al., 2023) | 14.5548 | 0.4271 | 0.6621 | 0.3706 | 0.2867 | 0.4024 | |
| COCO (Caesar et al., 2018) | TMPI (Khan et al., 2023) | 17.1759 | 0.5454 | 0.7314 | 0.3819 | 0.2028 | 0.2415 |
| 256x256 | AdaMPI (Han et al., 2022) | 23.6155 | 0.8263 | 0.9345 | 0.7232 | 0.1055 | 0.0797 |
| PLFNet+ (Ours) | 22.8933 | 0.7608 | 0.9070 | 0.5534 | 0.2378 | 0.2560 | |
| SinMPI (Pu et al., 2023) | 15.7272 | 0.4740 | 0.6712 | 0.3095 | 0.2392 | 0.3775 | |
| COCO (Caesar et al., 2018) | TMPI (Khan et al., 2023) | 18.1375 | 0.5444 | 0.7499 | 0.2906 | 0.1750 | 0.2758 |
| 512x512 | AdaMPI (Han et al., 2022) | 21.5931 | 0.7425 | 0.8925 | 0.5215 | 0.1156 | 0.1559 |
| PLFNet+ (Ours) | 20.2596 | 0.6229 | 0.8001 | 0.3007 | 0.2758 | 0.4226 |
@@
| PLFNet+ (Ours) | 0.0074 / 135 | 0.0086 / 116 |
@@ 5.1.1. Datasets
To evaluate the performance of PLFNet+ on view synthesis, we used two datasets. The first is a custom Blender dataset, created specifically for this task, which contains images from different scenes with camera movements constrained within a 20 cm cubical volume. The second dataset is COCO (Caesar et al., 2018), which we use during training in the same manner as AdaMPI (Han et al., 2022), where depth-image-based rendering (DIBR) with hole filling is employed to generate the training data. During evaluation, we take an input image and synthesize a novel image at a specified target location. The synthesized image is then treated as input to generate an image back at the original location. This process enables us to compare the reconstructed image with the ground truth, which is the original input image, allowing for an assessment of rendering quality. Both datasets are evaluated at resolutions of and and a movement within a volume of .
@@ 5.1.2. Evaluation Metrics
To objectively evaluate the quality of the synthesized novel views, we employed multiple standard and advanced metrics. These include peak signal-to-noise ratio (PSNR) and structural similarity index measure (SSIM) (Wang et al., 2004), as well as multi-scale structural similarity index measure (MS-SSIM) (Wang et al., 2003), pixel- based visual information fidelity (VIFP) (Sheikh and Bovik, 2006), deep image structure and texture similarity (DISTS) (Ding et al., 2020), and learned perceptual image patch similarity (LPIPS) (Zhang et al., 2018). These metrics capture not only pixel-level errors but also perceptual differences, providing a holistic view of the synthesis quality.
@@ 5.1.3. Results
Quantitative Results: The quantitative results for view synthesis across different datasets and resolutions are presented in Table 1. Our proposed PLFNet+ consistently outperforms AdaMPI (Han et al., 2022) on the Blender dataset and achieves competitive results on the COCO (Caesar et al., 2018) dataset. Although AdaMPI demonstrates slightly better performance in perceptual metrics such as DISTS and LPIPS, our method remains competitive across a range of metrics, especially on PSNR and SSIM.
Interestingly, we observe a slight performance drop in AdaMPI and PLFNet+ when increasing the resolution from to . This highlights the challenges of scaling models without increasing their capacity to handle higher-resolution inputs. In contrast to other methods that explicitly rely on depth estimation using the DPT model, our approach learns scene depth implicitly from RGB input alone, demonstrating the flexibility of our method without the need for pre-computed depth maps.
When evaluating inference time, PLFNet+ excels, particularly at higher resolutions. As shown in Table 2, our method achieves an inference time of 0.0086 seconds for a resolution of , outperforming all other methods by a significant margin. This makes PLFNet+ highly suitable for real-time applications, where rapid generation of novel views is critical, such as in dynamic video streams or interactive virtual environments.
Qualitative Results: Figure 4 provides a visual comparison of the view synthesis results for different methods on the COCO dataset. Each method’s output is accompanied by error maps, allowing for a more detailed analysis of their accuracy.
In contrast, both AdaMPI and our proposed PLFNet+ produce the most visually accurate results, with minimal artifacts. AdaMPI shows slightly better performance in some scenarios, particularly in handling fine details, but our method’s ability to efficiently handle reflections, occlusions, and fine textures stands out. While AdaMPI may have a slight edge in perceptual quality, our method’s superior inference speed makes it the more practical choice for real-time applications, balancing high-quality synthesis with the demands of time-sensitive environments.
5.2. Extending View Synthesis: Light Field Reconstruction
5.2.1. Datasets
We trained PLFNet on 75% of the Lytro Flowers and Stanford datasets, with the remaining 25% reserved for testing. The JPEG-Pleno dataset, however, was exclusively used for testing to evaluate the generalization capability of our model, as it was not included in the training process.
5.2.2. Evaluation Metrics
We utilize similar evaluation metrics as described in Section 5.1.2. Since the light field reconstruction works in very narrow baseline between each sub-aperture image, and has discrete camera locations, the search space is much smaller therefore we get much better scores in these metrics.
@@ 5.2.2. Evaluation Metrics
| PLFNet | 33.8982 | 0.9106 | 0.9838 | 0.8683 | 0.0590 | 0.0261 |
| Method | Purpose | Inference Time (s) |
| NoisyLFRecon | LF reconstruction | 1.1830 |
| DGLF | LF reconstruction | 1.5009 |
| DALF | LF reconstruction | 0.0583 |
| IR-V | LF reconstruction | 0.1429 |
| LFSphereNet | LF reconstruction | 0.0008 |
| PLFNet | Novel view synthesis | 0.0980 |
| NoisyLFRecon | DGLF | DALF | IR-V | LFSphereNet | PLFNet |

### v2

@@ 3.1. Positional Embedding
We aim to synthesize views from a single input image and a target pose , defined by 3D translation and rotation representing yaw, pitch, and roll, respectively. While recent approaches such as Cameras as Rays (Zhang et al., 2024) propose learning patch wise ray bundles for pose estimation, these are designed for multi-view image setups and rely on ray supervision and diffusion sampling. In contrast, our goal is real-time view synthesis from a single image, where camera pose is already known. Thus, we adopt a compact 6D vector representation and learn a high-dimensional embedding to support feed-forward synthesis at high frame rates.
Although direct regression using may suffice for small viewpoint changes, it becomes insufficient for larger baselines. To better capture pose dependent variations, we employ a positional embedding strategy that combines positional encoding, inspired by NeRF (Mildenhall et al., 2021), along with our proposed MLP module for enhanced representation of the target coordinate .
Coordinate Normalization: First, the target coordinates denoted by are normalized into to ensure consistency across varying ranges of viewpoints in scenes and training samples :
where denote the minimum and maximum values across the training set for each component of . This ensures all elements of the normalized pose vector . This normalization step ensures that the coordinates lie within a fixed range, which is crucial for subsequent encoding.
where is the normalized 6D coordinate vector, represents the encoding depth, and . The parameter is a scaling constant based on the domain of . For our experiments, we use , resulting in an output size of dimensions, which provides enhanced positional awareness (see Table 5 for ablation study results).
Learning Based Reprojection: Once the positional encoding has been computed, we further enhance the representation by passing it through an MLP. This MLP maps the given encoded vector into a much higher-dimensional space, which has proven to work well with scene regression task as shown in NeRF (Mildenhall et al., 2021). This process allows for a more expressive and flexible feature space which, unlike NeRF, can be adapted to different scenes since we only consider the coordinates in this representation and not the scene properties at this point. Our approach is similar to the adaptive layer normalization techniques used in NViST (Jang and Agapito, 2024).
Let be the output of the positional encoding, where is the dimensionality of the encoding. The MLP projects this vector into a high-dimensional space , where is chosen based on the desired resolution:
Then, the vector is reshaped into a feature map of size , where , w:h has the same aspect ratio as the original image, and . This reshaping operation can be expressed using the inverse vectorization notation as
Finally, the feature map is upsampled using nearest neighbor method to match the spatial resolution of the input image, denoted by , thus we set . This upsampling is performed using a scaling factor , as follows
@@ 3.2. Rendering Network
The role of the rendering network in PVSNet is to synthesize a target image from the input image and the positional embedding , obtained from the positional embedding stage. Our network architecture is inspired by LFSphereNet (Gond et al., 2023) and consists of two distinct encoder branches followed by a decoder network. These branches work in tandem to extract and merge spatial and positional features, ultimately generating high quality novel views.
@@ 3.2.2. Encoder II - Position-Aware Feature Extractor
The second encoder, , integrates both the input image and the positional embedding to adaptively capture position-specific features. Unlike the first encoder, which only processes the image, focuses on refining the feature maps based on the target position. This is done by capturing the high‐frequency variations in geometry and appearance much more effectively.
@@
where is the number of channels in . In our case, .
This stacked tensor is passed through a series of convolutional layers that preserve the spatial dimensions, ensuring pixel-level positional information is retained throughout the encoding process, as demonstrated by LFSphereNet (Gond et al., 2023). In the last two layers, Max Pooling is employed to downsample the feature maps to match the spatial dimension of feature map . The encoder produces a feature map , where is the number of feature channels. Additionally, skip connections are generated from intermediate layers for later use in the decoder.
@@ 3.3. Training Loss
To optimize our PVSNet, we combine four loss functions: L1 loss, Multi-Scale Structural Similarity (MS-SSIM) (Wang et al., 2003) loss, Focal Frequency Loss (FFL) (Jiang et al., 2021) and perceptual loss (Chen and Koltun, 2017) . Each of these losses contributes to different aspects of the synthesis process, improving both spatial and frequency domain accuracy.
@@
L1 Loss and MS-SSIM Loss are combined to capture fine-grained details, especially when dealing with larger baselines. The L1 loss minimizes pixel-wise differences between the predicted novel view image and the ground truth image , ensuring accurate reconstruction of intensity values. On the other hand, MS-SSIM is a perceptually motivated metric that focuses on preserving structural information and texture at multiple scales, correlating better with human visual perception, as demonstrated in image restoration tasks (Zhao et al., 2016).
@@
Focal Frequency Loss (FFL) is utilized to enhance the quality of the synthesized views, particularly in the frequency domain. We incorporate Focal Frequency Loss (FFL) (Jiang et al., 2021) to address spectral bias in neural networks by focusing on the frequency components that are hardest to synthesize. This helps the model to improve the reconstruction of high-frequency details, which are crucial for generating realistic novel views.
The FFL is defined as:
@@
where and represent the real and predicted frequency components at position in the Fourier domain, and is a dynamic weight that emphasizes the hardest-to-synthesize frequencies.
Perceptual Loss is used to improve the perceptual quality of the synthesized views by comparing high-level feature activations between the predicted and ground truth images. Inspired by (Chen and Koltun, 2017), we extract intermediate feature maps from a pretrained VGG-19 network and minimize the L1 difference between corresponding layers. These layers represent increasing levels of abstraction from edges to object parts and thus guide the synthesis model to preserve both fine grained detail and global structure. Specifically, we define the perceptual loss as:
where denotes the activation from layer of the VGG-19 network, and is a weighting factor inversely proportional to the number of elements in that layer. This perceptual supervision helps produce sharper, more semantically coherent outputs, particularly in areas where pixel-wise losses are insufficient to capture visual fidelity.
The total loss function for PVSNet is then defined as:
where is the weighting factor that balances the contribution of L1 and MS-SSIM losses and is a weight factor that balances the contribution of the FFL relative to the L1 and MS-SSIM losses. This combined loss function enables PVSNet to generate sharper and more realistic novel views by optimizing for both spatial accuracy and high-frequency details.
3.4. Implementation
The input of PVSNet is a single RGB image and the target position , for the positional embedding output we set since we use a single channel positional map. PVSNet was implemented in PyTorch111Code and checkpoint for our model will be made available upon acceptance of this paper at https://github.com/Realistic3D-MIUN/PVSNet , and trained with the following hyperparameters: batch size of 24, learning rate of 0.003, adam optimizer, MS-SSIM, L1 and FFL loss. After each 30 epochs, the learning rate was decreased by a factor of 0.2. The network was trained for 150 epochs on 2 GPUs (Nvidia A40) within a computing cluster with Intel Xeon Gold 6338 CPUs. The training duration was 48 hrs. Due to limitation of training data when using our synthetic dataset we employ multi stage training strategies to reach satisfactory results in the training process.
@@ 3.4.1. Multi Stage Training
Stage 1 serves as a preconditioning step, training the network on a large-scale light field dataset (Srinivasan et al., 2017). While the baseline between the views in each light field is relatively narrow, the large scale of dataset enables the network to learn both high and low level image features, implicitly understanding depth. In this stage, the network is conditioned to synthesize novel views within a constrained light field volume. Once the network demonstrates proficiency in this task, we transition to the next stage.
@@
Stage 3 involves training the network on our synthetic dataset (discussed in Section 4.1.1) containing random camera viewpoints within a cubical volume. The increase in the spatial search space necessitates the use of the positional embedding module for accurate view synthesis. In this stage, the network is fine-tuned to handle more complex camera translations and larger scene variations, fully utilizing the positional embedding to enhance rendering precision.
| Classroom | Italian Flat |
| Lone Monk | Barbershop |
@@ 4. Experimental Evaluation
In this section, we present a comprehensive evaluation of our proposed PVSNet architecture. We begin by assessing the performance of PVSNet for view synthesis on datasets with varying baselines and complexities. Following this, we demonstrate the capability of PVSNet for small baseline light field reconstruction and view synthesis within LF baseline from a single input image. The experiments for view synthesis are detailed in Section 4.1, while light field reconstruction results are discussed in Section 4.2.
@@
| PVSNet (Ours) | 31.2258 | 0.9386 | 0.9793 | 0.8702 | 0.0613 | 0.0275 | |
| PVSNet-Lite (Ours) | 28.7894 | 0.8902 | 0.9577 | 0.7655 | 0.0942 | 0.0566 | |
@@
| PVSNet (Ours) | 30.0085 | 0.9254 | 0.9733 | 0.8408 | 0.0733 | 0.0329 | |
| PVSNet-Lite (Ours) | 25.8741 | 0.8135 | 0.9100 | 0.6063 | 0.0852 | 0.0778 | |
| SinMPI (Pu et al., 2023) | 14.8544 | 0.4669 | 0.6484 | 0.3521 | 0.2940 | 0.5111 | |
| COCO (Caesar et al., 2018) | TMPI (Khan et al., 2023) | 17.6991 | 0.5822 | 0.7411 | 0.4032 | 0.2161 | 0.2490 |
| 256x256 | AdaMPI (Han et al., 2022) | 22.1027 | 0.8093 | 0.9241 | 0.7180 | 0.1152 | 0.0849 |
| PVSNet (Ours) | 24.0435 | 0.8462 | 0.9445 | 0.7518 | 0.0825 | 0.0600 | |
| PVSNet-Lite (Ours) | 21.3391 | 0.7555 | 0.9122 | 0.6195 | 0.1147 | 0.0915 | |
| SinMPI (Pu et al., 2023) | 15.1025 | 0.4561 | 0.6500 | 0.3149 | 0.2250 | 0.3555 | |
| COCO (Caesar et al., 2018) | TMPI (Khan et al., 2023) | 18.2453 | 0.6290 | 0.7851 | 0.3003 | 0.1750 | 0.2711 |
| 512x512 | AdaMPI (Han et al., 2022) | 21.3099 | 0.7356 | 0.8792 | 0.5099 | 0.1196 | 0.1673 |
| PVSNet (Ours) | 22.8850 | 0.7764 | 0.9015 | 0.6095 | 0.1070 | 0.1437 | |
| PVSNet-Lite (Ours) | 19.8910 | 0.6200 | 0.7842 | 0.4613 | 0.1364 | 0.1894 |
@@
| PVSNet (Ours) | 0.0161 / 62 | 0.0181 / 55 |
| PVSNet-Lite (Ours) | 0.0081 / 123 | 0.0096 / 104 |
4.1. View Synthesis
@@ 4.1.1. Datasets
To evaluate the performance of PVSNet on view synthesis, we use two datasets: a custom Blender dataset tailored for this task and the COCO dataset (Caesar et al., 2018) for more diverse scenes to demonstrate broader generalizability.
Blender dataset was created by selecting four freely available virtual scenes from the Blender demo website—namely Lone Monk, Classroom, Barbershop, and Italian Flat as shown in Figure 4. For each scene, we first positioned the camera at various initial locations and then applied random camera movements constrained within a cubical volume around each initial position. These random movements allowed us to capture multiple images per scene, all within the specified region. The resulting set of images from these scenes constitutes our custom Blender dataset222Dataset url will be made available upon acceptance of this paper.
COCO dataset was adopted with a similar approach to AdaMPI (Han et al., 2022) for generating training data, which involved utilizing depth-image-based rendering (DIBR) with hole filling. During evaluation, we take an input image and synthesize a novel image at a specified target location. The synthesized image is then treated as input to generate an image back at the original location. This process enables us to compare the reconstructed image with the ground truth (the original input image), thereby allowing for an assessment of rendering quality. Both datasets are evaluated at resolutions of and .
@@ 4.1.2. Evaluation Metrics
To objectively evaluate the quality of the synthesized views, we employed multiple standard and advanced metrics. These include peak signal-to-noise ratio (PSNR) and structural similarity index measure (SSIM) (Wang et al., 2004), as well as multi-scale structural similarity index measure (MS-SSIM) (Wang et al., 2003), pixel- based visual information fidelity (VIFP) (Sheikh and Bovik, 2006), deep image structure and texture similarity (DISTS) (Ding et al., 2020), and learned perceptual image patch similarity (LPIPS) (Zhang et al., 2018). These metrics capture not only pixel-level errors but also perceptual differences, providing a holistic view of the synthesis quality.
@@ 4.1.3. Results
In this section, we compare the view synthesis performance of PVSNet with state-of-the-art methods, namely AdaMPI (Han et al., 2022), TMPI (Khan et al., 2023), and SinMPI (Pu et al., 2023). To ensure a fair comparison, all methods were trained on the same dataset, using their respective original implementations and training strategies. Additionally, we introduce a lightweight variant called PVSNet-Lite, specifically optimized for maximum inference speed with minimal quality degradation. This version uses a simplified rendering network and reduced feature dimensionality to significantly boost FPS, making it ideal for latency-critical applications. The details of this PVSNet-Lite will be available in our open sourced code11footnotemark: 1.
Quantitative Image Quality: The quantitative results for view synthesis across different datasets and resolutions are presented in Table 1. Our proposed PVSNet consistently outperforms AdaMPI (Han et al., 2022) on the Blender dataset and achieves competitive results on the COCO (Caesar et al., 2018) dataset. Our method remains competitive across a range of metrics, especially on PSNR and SSIM.
Interestingly, we observe a slight performance drop in AdaMPI and PVSNet when increasing the resolution from to . This highlights the challenges of scaling models without increasing their capacity to handle higher-resolution inputs. In contrast to other methods that explicitly rely on depth estimation using the DPT (Ranftl et al., 2022) model, our approach learns scene depth implicitly from RGB input alone, demonstrating the flexibility of our method without the need for pre-computed depth maps.
Quantitative Processing Time: To better illustrate the trade off between inference speed and image quality, we present a comparison of all methods in Figure 5. The plot shows MS-SSIM versus FPS at both and resolutions, clearly highlighting that PVSNet achieves an optimal balance, while PVSNet-Lite pushes the boundary further by significantly improving FPS with only a modest drop in quality. As shown in Table 2, our method achieves an inference time of 0.0181 seconds (55 fps) for a resolution of , outperforming all other methods by a factor 18 . We also report results for PVSNet-Lite in Table 2, which achieves even higher frame rates (around 104 FPS at ) with only a marginal drop in visual quality. This makes PVSNet and PVSNet-Lite highly suitable for real-time applications, where rapid generation of novel views is critical, such as in dynamic video streams or interactive virtual environments. A pre-recorded video in our supplementary materials shows the performance in dynamic video streams.
Qualitative Image Quality: Figure 6 provides a visual comparison of the view synthesis results for different methods on the COCO dataset. Each method’s output is accompanied by error maps, allowing for a more detailed analysis of their accuracy.
@@
| PVSNet | 33.8982 | 0.9106 | 0.9838 | 0.8683 | 0.0590 | 0.0261 |
In contrast, both AdaMPI and our proposed PVSNet produce the most visually accurate results, with minimal artifacts. Our method’s ability to efficiently handle reflections, occlusions, and fine textures stands out, in addition the inference speed makes it the more practical choice for real-time applications, balancing high-quality synthesis with the demands of time-sensitive environments. Additional visual results when running the PVSNet on in-the-wild photos are presented in our supplementary materials video.
4.2. Light Field Reconstruction and View Synthesis
In this section, we demonstrate alternative approach for view synthesis where Light Field reconstruction is the intermediate step. We focus on capability of proposed PVSNet architecture in reconstructing full light fields from a single input image. In this case study, we target very small baseline of the LFs, and we use a smaller version of PVSNet by modifying the positional embedding step to only use the coordinate normalization during training and inference. The setup of PVSNet which utilizes positional embedding with only the coordinate normalization has been studied in detail in our ablation study in Section 5.1, where we show that this setup is sufficient to handle small baselines.
4.2.1. Datasets
We trained PVSNet on 75% of the Lytro Flowers and Stanford datasets, with the remaining 25% reserved for testing. The JPEG-Pleno dataset, however, was exclusively used for testing to evaluate the generalization capability of our model, as it was not included in the training process.
4.2.2. Evaluation Metrics
We utilize similar evaluation metrics as described in Section 4.1.2. Since the light field reconstruction works in very narrow baseline between each sub-aperture image, and has discrete camera locations, the search space is much smaller. Therefore we expect much better scores in these metrics.
| Method | Purpose | LF Reconstruction (sec) | View Synthesis (sec) |
| NoisyLFRecon | LF reconstruction | 1.0923 | 2.3933 |
| DGLF | LF reconstruction | 1.3062 | 2.6072 |
| DALF | LF reconstruction | 0.0577 | 1.3587 |
| IR-V | LF reconstruction | 0.1342 | 1.4352 |
| LFSphereNet | LF reconstruction | 0.0008 | 1.3018 |
| PVSNet | View synthesis | 0.0850 | 0.0079 |
| NoisyLFRecon | DGLF | DALF | IR-V | LFSphereNet | PVSNet |